OpenStudy (anonymous):

The lengths of the legs of a right triangle are 12 cm and 16 cm. What is the length of the hypotenuse? 4 centimeters 20 centimeters 22 centimeters 28 centimeters

OpenStudy (anonymous):

I need alot of help for this please

OpenStudy (butterfield1215):

go onto mathway.com

OpenStudy (anonymous):

You can use the Pythagorean Theorem

OpenStudy (butterfield1215):

you should get B

OpenStudy (anonymous):

I did thanks bro

OpenStudy (butterfield1215):

medal pls

OpenStudy (anonymous):

yup

OpenStudy (vhtran):

pythagorean theorem states that a\[a ^{2} + b ^{2}=c ^{2}\] If \[a ^{2}\] is a leg and \[b ^{2}\] is the other leg, you can solve for c\[c^{2}\ So the next time you solve for this question you should use pythagorean theorem. Plug in A as the first leg and B as the other. Once you have squared them both add them together and then find the square root of it. Do this to all other questions you have involving 2 legs and finding the last side length. Happy to help ^u^

OpenStudy (anonymous):

20

OpenStudy (er.mohd.amir):

in right angle triangle hypotenuse^2=(side1)^2+(side2)^2 so H*H=12*12+16+16=144+256=400=20*20 H=20